OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 65 miles per hour (mph), and the other car is traveling 50 mph. How long will it take before the cars are 575 miles apart?

OpenStudy (anonymous):

Separation rate is 50 + 65= 115 mph. Recall that d/r=t, so 575/115=5 hours.

OpenStudy (vishal_kothari):

65(T) + 50(T) = 575 5 hrs